1. (x2-4xy-2y2)dx + (y2-4xy-2x2)dy=0
  2. (x2-xtan2y+sec2y)dy = (tany-2xy-y)dx
  3. (2xy4ey+2xy3+y)dx + (x2y4ey-x2y2-3x)dy=0
  4. (y/x secy-tany)dx+(secy logx-x)dy=0
  5. y dy/dx +4x/3 -y2/3x =0
1
Expert's answer
2021-06-01T03:34:54-0400

Solutions:-

1.

"P(x,y)=x^2-4xy-2y^2"

"Q(x,y)=y^2--4xy-2x^2"

"\frac{\partial P(x,y)}{\partial y}=\frac{\partial Q(x,y)}{\partial x}=-4x-4y"

"\frac{\partial f(x,y)}{\partial x}=P(x,y)"

"\frac{\partial f(x,y)}{\partial y}=Q(x,y)"


"f(x,y)=\int(x^2-4xy-2y^2)dx=x^3/3-2x^2y-2xy^2+g(y)"

"\frac{\partial f(x,y)}{\partial y}=\frac{\partial }{\partial y}(x^3/3-2x^2y-2xy^2+g(y))=-2x^2-4xy+\frac{dg(y)}{dy}"

"-2x^2-4xy+\frac{dg(y)}{dy}=-2x^2-4xy+y^2"

"\frac{dg(y)}{dy}=y^2"

"g(y)=y^3\/3"

"f(x,y)=x^3\/3-2x^2y-2xy^2+y^3\/3"


The solution is "f(x,y)=c"

"x^3\/3-2x^2y-2xy^2+y^3\/3=c"


2.

"P(x,y)=y+2xy-tany"

"Q(x,y)=x^2--sec^2y-xtan^2y"

"\frac{\partial P(x,y)}{\partial y}=\frac{\partial Q(x,y)}{\partial x}=2x-sec^2y+1"

"\frac{\partial f(x,y)}{\partial x}=P(x,y)"

"\frac{\partial f(x,y)}{\partial y}=Q(x,y)"


"f(x,y)=\int(y+2xy-tany)dx=xy+x^2y-xtany+g(y)"

"\frac{\partial f(x,y)}{\partial y}=\frac{\partial }{\partial y}(xy+x^2y-xtany+g(y))=x+x^2-xsec^2y+\frac{dg(y)}{dy}"

"x+x^2-xsec^2y+\frac{dg(y)}{dy}=x^2--sec^2y-xtan^2y"

"\frac{dg(y)}{dy}=-sec^2y"

"g(y)=-tany"


"f(x,y)=xy+x^2y-tany-xtany"

"xy+x^2y-tany-xtany=c"


5.

"\frac{3yy'}{y^2-4}=x"


"\int\frac{3ydy}{y^2-4}=\int xdx"


"3ln(y^2-4)=x^2+c_1"

"y=\sqrt{ce^{x^2/3}+4}"


3.

"Mdx+Ndy=0"

"M=2xy^4e^y+2xy^3+y"

"N=x^2y^4e^y-x^2y^2-3x"

"\frac{\partial M}{\partial y}=2x(y^4e^y+4y^3e^y)+6xy^2+1"

"\frac{\partial N}{\partial x}=2xy^4e^y-2xy^2-3"


"\frac{\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x}}{M}=\frac{-8xy^2-4-8xy^3e^y}{2xy^4e^y+2xy^3+y}=-\frac{4}{y}"


Integrating factor:

"e^{\int(-4/y)dy}=1/y^4"

Multiplying by the I.f. we get:

"(2xe^y+2x\/y+1\/y^3)dx+(x^2e^y-x^2y^2-3x\/y^4)dy=0"

which is exact.

Then:

"\int (2xe^y+2x/y+1/y^3)dx=x^2e^y+x^2/y+x/y^3"


Solution:

"x^2e^y+x^2\/y+x\/y^3=c"


4.

"\frac{\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x}}{M}=-tany"

I.f.:

"e^{\int-tanydy}=cosy"


Then:

"(y\/x secy-tany)cosydx+(secy logx-x)cosydy=0"

"(y\/x-sinx)dx-(xcosy-logx)dy=0"

"\int (y/x-sinx)dx=ylogx-xsiny"


Solution:

"ylogx-xsiny=c"